OpenStudy (anonymous):

Given log2(x), show in terms of ln(x) Hint: set log2(x) = w

OpenStudy (anonymous):

clarification \[\log_{2}X \]and \[\ln x\]

OpenStudy (anonymous):

\[\log_{b}a = \frac{\log_{c} a}{\log_{c} b}\]for any base, c. Also,\[\ln x = \log_{e}x\]

OpenStudy (anonymous):

But maybe they don't want you to use that shortcut.\[\log_{2}x = w\]implies\[2^{w} = x\]